OpenStudy (anonymous):

Okay somebody please help me figure out if this piecewise function is right and how to graph it? (Will give medal & fan) I had to make a function for this scenario: "A local store owner pays her employees time-and-a-half for overtime. That means if an employee works more than 40 hours per week, the store will pay him 1.5 times his regular hourly wage of $6.50." "Write and graph a piecewise function that illustrates an employee's weekly pay, W, as a function of the number of hours, h, that employee works"

OpenStudy (anonymous):

The function I have is f(h) = 6.50*40 + (h - 40)9.75 or f(h) = 260 + (h - 40)9.75, I'm not quite sure where the W goes though...

OpenStudy (anonymous):

@phi I'm sorry I keep asking you for help but you usually seem of more help to me than most others /:

OpenStudy (anonymous):

hey @TammisaurusRex i think the w goes for the weekly pay so i assume it goes after 40

OpenStudy (anonymous):

so, f(h) = 260 + (h - 40w)9.75?

OpenStudy (phi):

W, as a function of the number of hours, h, means W(h)= stuff in other words, instead of f(h), just use W(h)

OpenStudy (anonymous):

so itwould be 40w yes

OpenStudy (anonymous):

ohhh!! okay @phi , that makes sense. so W(h) = 260 + (h - 40)9.75?

OpenStudy (anonymous):

And how would I graph that?

OpenStudy (anonymous):

sorry i couldnt help

OpenStudy (phi):

yes, but that is only true if h≥40 hours if they work 20 hours, you will get the wrong answer. that is why they say Write and graph a *piecewise* function you will have a function with two different pieces

OpenStudy (anonymous):

Oh okay.. how do I do that?

OpenStudy (phi):

you write down W(h)= whatever works for when h<40 and then write W(h) = equation 1 if h<40 = equation 2 if h≥ 40

OpenStudy (phi):

See http://www.mathsisfun.com/sets/functions-piecewise.html and scroll down to And this is how you write it: to see how one of these looks

OpenStudy (anonymous):

so W(h) = 6.50h if h<40 W(h) = 260 + (h - 40)9.75 if h≥40 ?

OpenStudy (anonymous):

oh okay hold on reading that now

OpenStudy (anonymous):

Oh okay I see

OpenStudy (phi):

yes, though people use W(h) = { and then the two equations (see the link for the example) But you can write it the way you did if you can't do it the "official way"

OpenStudy (anonymous):

It would be a bit hard for me to do it the "official way" on microsoft word, lol so I'll write it the way I have it.

OpenStudy (anonymous):

Thank you!

OpenStudy (phi):

microsoft word has an equation editor that would let you do it that way...

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

Oh really? How do I find it?

OpenStudy (anonymous):

And @C.A.M it's fine

OpenStudy (anonymous):

Oh okay, I got it. Thank you (:

OpenStudy (phi):

Of course you have to learn how to use the equation editor. You could put that off for another day if you don't have the time. But it does make for nicer write-ups.

OpenStudy (phi):

Here is what your function looks like when graphed